OpenStudy (anonymous):

hello how do I do this: 8w-2w(to the 2nd power, computer wont let me input power)-6 divided by w-1

OpenStudy (anonymous):

Is the whole expression 8w - 2w to be squared? Or is it just the w?

OpenStudy (anonymous):

If it is (8w - 2w)^2, then (8 - 2w)^2 = (6w)^2 = 36w^2

OpenStudy (anonymous):

hi sorry just the w

OpenStudy (anonymous):

Actually, it appears you are asking what 8w - 2w^2 -6 divided by w - 1 is. -2w + 6 ________________ w - 1 | -2w^2 + 8w - 6 -( -2w^2 +2w) ____________ 6w - 6 -(6w - 6) ________ 0

OpenStudy (anonymous):

Alternatively, -2w^2 + 8w - 6 = -2(w^2 -4w +3) = -2(w - 3)(w - 1) If you multiply the -2 by w - 3, you get -2w + 6, the same as we got by long division.

OpenStudy (anonymous):

Is this the actual problem you need help with, or did I misread your question?

OpenStudy (anonymous):

the answer was -2w+10+ 4/w-1 using long division according to another site. I will rework my long division to ck.. thnks